# The Learning Approach to Games

### v2

@@ 2 Discrete Games
Let $\mathbb{T}=\mathbb{N}$ denote the time indices, define $\mathbb{T}_{s}^{t}:=\{t,\dots,s\}$ for each $t\leq s\in\mathbb{T}$, and set $\mathbb{T}_{t}:=\mathbb{T}_{t}^{0},\mathbb{T}^{t}:=\mathbb{T}_{\infty}^{t}$. Let $\mathbb{S}_{t}$ be a measurable state space for each $t\in\mathbb{T}$, and define $\mathbb{S}:=\bigcup_{t\in\mathbb{T}}\mathbb{S}_{t}$. For arbitrary set $E$ with Borel $\sigma$-algebra ${\cal B}(E)$, let ${\cal P}(E)$ denote the set of all probability measures on $E$. We will always consider discrete indexing to avoid discussions on regularities and measurability.
Take $\Omega:=\prod_{t\in\mathbb{T}}\mathbb{S}_{t}$ as the canonical space. Define $X:\mathbb{T}\times\Omega\rightarrow\mathbb{S}$ as the canonical process: $X_{t}:\Omega\to\mathbb{S}_{t}$, and $X_{t}(\omega)=\omega_{t}$ for each $\omega\in\mathbb{X}$ and $t\in\mathbb{T}$. Let $\mathbb{F}^{X}$ denote the filtration generated by $X$. We always require any function defined on $\mathbb{T}\times\Omega$ to be Markovian, similar to the canonical process, and denote their parameters as $(t,x)$ where it is understood that $x\in\mathbb{S}_{t}$. Note that since $\mathbb{S}_{t}$ is also indexed, this does not impose a restriction. In particular, there is no need to bookkeep paths of the process.
We use $i\in\mathbb{N}_{0}:=\mathbb{N}\setminus\{0\}$ as the index for players. Naturally, it is common to restrict models to finitely many players, but it is not necessary for our purposes. For any $i\in\mathbb{N}_{0}$, let $\mathbb{A}^{t,x;i}$ be the action space of player $i$ at $(t,x)\in\mathbb{T}\times\mathbb{S}$. Introduce
@@
As for the learning parameters, we will now begin to introduce horizon, transitions between states, transition costs, state values, potential behaviors of other players, optimal controls and expectations of the players. Our choices are inherently limited as a player might be arbitrarily complicated, and includes some modeling choices, but general enough to cover a diverse set of examples and to demonstrate the concept of uncertain equilibrium. It is important to note that our primary goal is not to model the learning of these parameters, as each parameter could constitute its own line of research. For simplicity, we will temporarily disregard the index $i$ and focus solely on the perspective of a single player.
First of all, players need to have a horizon $\hat{T}$. We assume that, as always the case in reality, players cannot predict the future indefinitely with reasonable accuracy. In other words, as the horizon of prediction increases, the distribution of the state process contains progressively less useful information, eventually rendering it useless. Even under the infinite horizon settings, typically one introduces a discounting factor, essentially approximating a finite horizon. Thus, let
@@
Next, the player might have an estimate of the transition probabilities;
@@
$\hat{p}(t,x,\vec{a};\cdot)\ \text{ is a probability measure on }\ \mathbb{S}_{t+1},\ \ \text{for all}\ t\in\mathbb{T},x\in\mathbb{S}_{t},\text{ and }\vec{a}\in\vec{\mathbb{A}}^{t,x}$
Similarly, introduce $\mathscr{M}_{p}$ as the space of all such mappings in (2.2). Note that Large Language Models excel in the task of constructing $\hat{p}$ and solve the universal problem of predicting the next input in the specific context of language.
Given $\hat{p}$ as in (2.2), an initial $(t,x)\in(\mathbb{T},\mathbb{S}_{t})$, and $\vec{\alpha}\in\vec{\cal A}$, player induces a distribution $\mathbb{P}^{t,x,\vec{\alpha}}:=\mathbb{P}^{\hat{p};t,x,\vec{\alpha}}$ for the canonical process as usual; for all $t\leq s$ and $(\tilde{x},y)\in(\mathbb{S}_{s},\mathbb{S}_{s+1})$,
$\mathbb{P}^{t,x,\vec{\alpha}}(X_{s+1}=y|X_{s}=\tilde{x})=\hat{p}(s,\tilde{x},\vec{\alpha}(s,\tilde{x});y),\ \ \text{and}\ \ \mathbb{P}^{t,x,\vec{\alpha}}(X_{t}=x)=1$ (2.3)
@@
    Here, we assume that players model the others potential controls depending on their own control. However, one might model that this depends on the path of states of the players, or in fact, any other observables are legitimate as long as the cost (2.6) is well-defined. Our choice here is to illustrate the optimization in a visible manner.
  • (ii)
@@
    It is crucial for players to learn reliable $\hat{\Gamma}$. To not only compete with but also cooperate with other players, they must generate reliable estimates of the actions of others. We will show in the two-player game discussed in Section 2.1 that because the costs to the players depend on each other’s states, omitting this aspect from the player model won’t accurately capture the observed dynamics.
A crucial notion to introduce is the value of a player. We observe that value is fundamentally a formalism used to determine an optimal action. We do not regard it as something tangible that a player necessarily obtains, which might only be the case in limited situations. However, even in these situations, value is typically considered an expectation rather than a definitive quantity. Now, we introduce two concepts: transition costs and state values, which are akin to running costs and terminal costs in the literature [^3] We use cost and value interchangeably. In the case of scalar objectives as in this work, distinction is more pronounced. However, for multi-objective frameworks, there is typically no binary choice, but rather a continuum of choices.:
$\hat{F}:\hat{\Omega}\times\mathbb{T}\times\Omega\times\vec{\mathbb{A}}\to\mathbb{R},\qquad\hat{\phi}:\hat{\Omega}\times\mathbb{T}\times\Omega\to\mathbb{R},$ (2.5)
and let $\mathscr{M}_{F},\mathscr{M}_{\phi}$ denote the sets of mappings as in (2.5). An important difference is that the player models these as random variables on some probability space $(\hat{\Omega},\hat{\cal F},\hat{\mathbb{P}})$. In particular cases, it might be useful to characterize the measure space $(\hat{\Omega},\hat{\cal F},\hat{\mathbb{P}})$, however, once can also fix the sufficiently large probability space $([0,1],{\cal B}([0,1]),\textbf{m})$ and concentrate on the random variables. We remark that state value $\hat{\phi}$, in particular, induces an ordering on states. Moreover, reaching a particular state by different intermediate paths, or different set of actions might have varying costs, which is aimed to be captured by the transition cost $\hat{F}$.
Now, given $(\hat{T},\hat{p},\hat{\Gamma},\hat{F},\hat{\phi})$, the value of player becomes
$J(t,x;\alpha):=\int_{\vec{\cal A}}J(t,x;\vec{\alpha})\hat{\Gamma}_{t}(\alpha;d\vec{\alpha})\ \ \text{where}$ (2.6)
$J(t,x;\vec{\alpha}):=\mathbb{E}^{t,x,\vec{\alpha}}\Big{[}\hat{\phi}(t+\hat{T},X_{t+\hat{T}})+\sum_{s=t}^{t+\hat{T}-1}\hat{F}(s,X_{s},\vec{\alpha}(s,X_{s}))\Big{]},\quad\mathbb{E}^{t,x,\vec{\alpha}}:=\mathbb{E}^{\mathbb{P}^{t,x,\vec{\alpha}}}$
@@
which is a random variable on $\hat{\Omega}$. Set $\mathscr{M}_{J}^{i}$ as the space of all such functions $(\hat{\Omega}\times\mathbb{T}\times\Omega\times{\cal A}^{i}\to\mathbb{R})$. We point out that there is no fundamental difference between requiring either a distribution or a random variable, but the latter allows us to refer to samples.
Let us recall the game of chess, which serves as an excellent example to keep in mind throughout this work. In chess, $\hat{p}$ yields deterministic transitions. However, a player does not know what actions the opponent will take within $\{t,\dots,t+\hat{T}\}$, and beyond that, it is unclear what the transition costs of actions or the value of being in a particular state at $t+\hat{T}$ might be. These are all crucial components for a player to learn. Notably, the heuristic values assigned to pieces are designed to guide players in learning $\hat{F}$ and $\hat{\phi}$. While simplistic, these heuristics serve as an initial guide. Moreover, as we have mentioned, knowledge about the opponent can alter the values of strategies, which is captured in (2.6). Let us also emphasize that the player’s horizon may depend significantly on the current state. Towards the endgame, for instance, a well-trained chess player might be able to estimate many steps ahead, whereas this ability may be considerably more limited during the middle stages of the game.
We remark that it is obviously equivalent to define
$J(t,x;\vec{\alpha}):=\mathbb{E}^{t,x,\vec{\alpha}}\Big{[}\hat{\phi}(t+\hat{T},X_{t+\hat{T}})-\hat{\phi}(t,x)+\sum_{s=t}^{t+\hat{T}-1}\hat{F}(s,X_{s},\vec{\alpha}(s,X_{s}))\Big{]}$
@@
which is more intuitive and useful for computations, but we will not keep track of it. Moreover, one can further extend the space of parameters. A simple example would be to write $\hat{F}(t,s,x,a)$. We are merely pointing out basic modeling choices, with the understanding that these will need to be tailored to the specific problem at hand. In fact, a truly complex player, such as a human, would not only adapt existing estimates and behaviors, but also develop new types of estimates and behaviors over time. Just as a child starts with limited capacities and gradually acquires a rich repertoire, such dynamic structural growth might indeed be essential for modeling higher-order intelligence. While this remains far beyond current work, it represents a compelling direction for future research.
@@
And then, to incorporate the potential difficulty and uncertainty in identifying the optimal control, we introduce the next learning parameter;
$\hat{\pi}:\hat{\Omega}\times\mathbb{T}\times\Omega\to{\cal P}({\cal A}^{i})\qquad\text{where},$ (2.7)
$\hat{\pi}(\hat{\omega},t,x)(d\alpha)=\hat{\pi}(\hat{\omega},t,x)(d\tilde{\alpha})\ \ \text{whenever}\ \ \alpha=^{t,x;i}\tilde{\alpha}$
Here, at $(\hat{\omega},t,x)$, $\hat{\pi}$ approximates the potential optimal controls for $J(\hat{\omega},t,x,\cdot)$, which will be dictated by the equilibrium condition below. Note that, the equality should be understood in terms of two subsets of controls having equal probabilities, when they are equal once extended with the relation $=^{t,x;i}$. Also, we didn’t suppress $\hat{\omega}$ as it will be integrated.
@@
Now, the crucial observation is that, even when optimal control can be solved exactly, uncertainty over the value will naturally induce a probability distribution over controls. That is, given $(\hat{T},\hat{p},\hat{\Gamma},\hat{F},\hat{\phi},\hat{\pi})$, we define ${\Upsilon}^{t,x;i}\in{\cal P}({\cal A}^{t,x;i})$ as
${\Upsilon}^{t,x;i}(d\alpha):=\int_{\hat{\Omega}}\hat{\pi}(\hat{\omega},t,x)(d\alpha)\hat{\mathbb{P}}(d\hat{\omega})$ (2.8)
We have introduced distributions over controls, which then further induces distribution over actions:
$\gamma^{t,x;i}(da):={\Upsilon}^{t,x;i}\big{(}\alpha:\ \alpha(t,x)=da\big{)}\in{\cal P}(\mathbb{A}^{t,x;i})$ (2.9)
Although apriori ${\Upsilon}^{t,x;i}$ appears to be solely induced by $\hat{\pi}$, $\hat{\pi}$ itself is a function of the value $J^{i}$, and hence ${\Upsilon}^{t,x;i}$ is a function from $\mathscr{M}_{T}\times\mathscr{M}_{p}\times\mathscr{M}_{\Gamma}\times\mathscr{M}_{F}\times\mathscr{M}_{\phi}\times\mathscr{M}_{\pi}$. We introduced $\hat{\pi}$ to carry out the discussions more explicitly. [^4] When $\hat{\pi}$ is a determined function of the underlying estimates, as in the case of our examples in this work, it is more appropriate to embed it into the behavior ${\Upsilon}$. In the case where strategies are learned both from observations and other estimates, it is even more favorable to adjust the Definition 1 slightly so that the domains can include spaces of estimations. We avoid this, as it would not enhance the discussion but would complicate the notation. In the two-player example discussed in section 2.1, $\hat{\pi}$ solves the optimization by brute force, yielding a deterministic action for each scenario in $\hat{\Omega}$. In the control problem described in section 4, $\hat{\pi}$ constructs the distribution over strategies relying again on the associated values. Let us note that we induce a distribution over controls because it is a more familiar and convenient choice; however, by considering sequences of future states, one can easily induce a distribution over sequences of actions.
@@
This behavior, represented by $\hat{\pi}$, is well suited for the learning or playing phase and is, in essence, Thompson sampling adapted to our framework. During a competition phase, one might modify $\hat{\pi}$ to be deterministic, selecting the control corresponding to the mode of the value. In particular situations, such as performing surgery, it is not only wrong but also unethical to forgo the most likely action and instead select one at random. Likewise, a deterministic $\hat{\pi}$ might instead choose the control that maximises a confidence bound of the value. Moreover, a player may adjust its behavior to occasionally select unlikely actions, exploring states that are disadvantageous or even entirely unseen when facing a weak opponent.
@@
It is important to motivate the role of randomness in value, which then induces a probability distribution over actions in (2.8). Recall that in sufficiently complex settings, such as chess, values are inherently unknown and must be learned through significant effort. That is, the randomness of the value models what is unknown to the player. The key role of randomness in value is to allow players to explore systematically. If the player is not satisfied with the current estimates, it is natural to assign greater probabilities to unexplored controls for having higher values. This approach naturally leads to the search for controls with more satisfactory results. We will demonstrate a toy version for the two-player game in Section 2.1. Consider a daily life scenario: suppose you have a favorite dish, and your friend suggests a new one you have never tried. Since you expect it is not better, if your model only considers expectations, you won’t explore. However, if you model the value as random, you might try the new dish. This approach aligns with the common intuition that a better understanding of values should lead to less uncertain strategies. Once you try the new dish, you know its value to you more accurately, and that experience might deter further exploration.
Now that we have introduced the spaces of estimations and behaviors, let us also recall the observations. These observations might be coming from real world experience. Or in the mean-field regime, players can generate observations by assuming every other player is exactly the same. To distinguish the index of observations and learning from time, we reserve $n\in\mathbb{N}$ written on the left superscripts. Let ${\cal E}^{i}$ denote the space of observables for player $i$, and set $\mathscr{E}^{i}$ as the finite sequences of ${\cal E}^{i}$. Similar to previous notations, set $\vec{\cal E}:=\prod_{i\in\mathbb{N}_{0}}{\cal E}^{i}$ and $\vec{\mathscr{E}}:=\prod_{i\in\mathbb{N}_{0}}\mathscr{E}^{i}$. Now, observations are given as
${\cal O}^{i}:\Omega^{u}\times\mathbb{N}\to\mathscr{E}^{i}\ \ \text{where}\ \ {}^{n}{\cal O}^{i}\ \ \text{is a subsequence of}\ \ {}^{n+1}{\cal O}^{i},\ \ \forall i\in\mathbb{N}_{0},\ n\in\mathbb{N},\ w^{u}\in\Omega^{u}$ (2.10)
and set $\vec{\cal O}=({\cal O}^{1},{\cal O}^{2},\dots)$. Recall that we assume there exists a universal probability distribution $\mathbb{P}^{u}$ on some arbitrary $\Omega^{u}$, which might be induced by real dynamics of the universe that players are in and distributions players are drawing from. We say that ${}^{n}{\cal O}^{i}$ characterizes all the observations of player $i$ up to age $n$.
@@
Now, we want to acknowledge the existence of a learning algorithm. We say a collection of functions $\mathfrak{L}_{\varphi}^{i}$ for $\varphi\in\{T,p,\Gamma,F,\phi,\pi\}$ is the learning algorithm of player $i$. Recall that $\mathscr{M}_{T},\mathscr{M}_{p},\mathscr{M}_{\Gamma},\mathscr{M}_{F},\mathscr{M}_{\phi},\mathscr{M}_{\pi}$ are the spaces of estimations as in (2.1), (2.2), (2.4), (2.5), and (2.7) respectively. Then, we let
@@
$\mathfrak{L}_{\varphi}^{i}:\mathscr{E}^{i}\to\mathscr{M}_{\varphi},\quad\forall\varphi\in\{T,p,\Gamma,F,\phi,\pi\}$ (2.11)
Moreover, for any function $\mathfrak{L}_{\varphi}^{i}$ on $\mathscr{E}^{i}$, we set estimations of the player as
@@
${}^{n}\hat{\varphi}^{i}:={}^{n}\mathfrak{L}_{\varphi}^{i}:=\mathfrak{L}_{\varphi}^{i}({}^{n}{\cal O}^{i}),\quad\forall\varphi\in\{T,p,\Gamma,F,\phi,\pi\}$
and call ${}^{0}\hat{\varphi}^{i}$ as the prior of player $i$. Note that, $({}^{n}\hat{T}^{i},{}^{n}\hat{p}^{i},{}^{n}\hat{\Gamma}^{i},{}^{n}\hat{F}^{i},{}^{n}\hat{\phi}^{i})$ defines ${}^{n}J_{\mathfrak{L}}^{i}(t,x;\alpha)$ as in (2.6), which then together with ${}^{n}\hat{\pi}^{i}$ induces ${\Upsilon}_{\mathfrak{L}}^{t,x;i}$ on ${}^{n}{\cal A}^{t,x;i}$ as in (2.8) and ${}^{n}\gamma_{\mathfrak{L}}^{t,x;i}$ on $\mathbb{A}^{t,x;i}$ as in (2.9). [^5] We will drop $\mathfrak{L}$ from subscripts if it is clear from the context.
We say all players are behaving rationally, if at each $(t,x)\in\mathbb{T}\times\mathbb{S}$, they draw an action from the distribution ${}^{n}\vec{\gamma}_{\mathfrak{L}}^{t,x}:=({}^{n}\gamma_{\mathfrak{L}}^{t,x;1},{}^{n}\gamma_{\mathfrak{L}}^{t,x;2},\dots)$, or draw from the distribution ${}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x}:=({}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x;1},{}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x;2},\dots)$ to decide on strategies for $\mathbb{T}_{t+\hat{T}}^{t}$. [^6] These are equivalent under time-consistency. The concept is discussed later in this section. We always assume that players are learning under rational behaviors. In a two-player game, for example, if one player behaves nonsensically, expecting any type of convergent behavior becomes meaningless. We emphasize that this assumption of rationality is not a constraint but rather means that we model players within the current framework of this work.
For notational convenience, instead of explicitly denoting $\{T,p,\Gamma,F,\phi,\pi\}$, we keep using $\varphi$ to indicate any one or all of the estimations, and denote $\vec{\varphi}:=(\varphi^{1},\varphi^{2},\cdots)$.
@@
We say $\vec{\varphi}\in\mathscr{M}_{\varphi}^{\mathbb{N}_{0}}$ is an $(\varepsilon,r,\delta)$-uncertain equilibrium at $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$ under the learning algorithm $\vec{\mathfrak{L}}_{\varphi}$ if,

### v4

@@ 3 Discrete Games
Let $\mathbb{T}=\mathbb{N}$ denote the time indices, and $\mathbb{S}_{t}$ be a measurable state space for each $t\in\mathbb{T}$. Set $\mathbb{S}:=\bigcup_{t\in\mathbb{T}}\mathbb{S}_{t}$. For arbitrary set $E$ with Borel $\sigma$-algebra ${\cal B}(E)$, let ${\cal P}(E)$ denote the set of all probability measures on $E$. We will always consider discrete indexing to avoid discussions on regularities and measurability.
Take $\Omega:=\prod_{t\in\mathbb{T}}\mathbb{S}_{t}$ as the canonical space. Define $X:\mathbb{T}\times\Omega\rightarrow\mathbb{S}$ as the canonical process: $X_{t}:\Omega\to\mathbb{S}_{t}$, and $X_{t}(\omega)=\omega_{t}$ for each $\omega\in\mathbb{X}$ and $t\in\mathbb{T}$. Let $\mathbb{F}^{X}$ denote the filtration generated by $X$. We always require any function defined on $\mathbb{T}\times\Omega$ to be Markovian, similar to the canonical process, and denote their parameters as $(t,x)$ where it is understood that $x\in\mathbb{S}_{t}$.
We use $i\in\mathbb{N}_{0}:=\mathbb{N}\setminus\{0\}$ as the index for players. For any $i\in\mathbb{N}_{0}$, let $\mathbb{A}^{t,x;i}$ be the action space of player $i$ at $(t,x)\in\mathbb{T}\times\mathbb{S}$. Introduce
@@
For the connection between players and the environment, let ${\cal E}^{i}$ denote the space of observables for player $i$, and set $\mathscr{E}^{i}$ as the finite sequences of ${\cal E}^{i}$. Similar to previous notations, set $\vec{\cal E}:=\prod_{i\in\mathbb{N}_{0}}{\cal E}^{i}$ and $\vec{\mathscr{E}}:=\prod_{i\in\mathbb{N}_{0}}\mathscr{E}^{i}$. We will state every estimate as depending on $(t,x)$, and hence we will assume $\mathbb{T}\times\Omega\subset{\cal E}^{i}$. Also, set $\mathscr{A}^{i}$ as finite sequences in $\mathbb{A}^{i}$, and $\vec{\mathscr{A}}:=\prod_{i\in\mathbb{N}_{0}}\mathscr{A}^{i}$.
As for the learning parameters, we will now begin to introduce horizon, transitions between states, transition costs, state values, potential behaviors of other players, optimal controls and expectations of the players. Our choices are inherently limited as a player might be arbitrarily complicated. Our aim here is to restate the general setting for many of our games in the perspective of player , and demonstrate the concept of uncertain equilibrium. For simplicity, we will temporarily disregard the index $i$ and focus solely on the perspective of a single player.
First, we let players have a horizon $\hat{T}$. Players cannot predict the future indefinitely with reasonable accuracy. In other words, as the horizon of prediction increases, the distribution of the state process contains progressively less useful information, eventually rendering it useless. Thus, let
@@
Next, the player have an estimate of the transition probabilities;
@@
$\hat{p}(t,x,\vec{a};\cdot)\ \text{ is a probability measure on }\ \mathbb{S}_{t+1},$
$\text{for all}\ t\in\mathbb{T},x\in\mathbb{S}_{t},\text{ and }\vec{a}\in\vec{\mathbb{A}}^{t,x}.$
Similarly, introduce $\mathscr{M}_{p}$ as the space of all such mappings in (3.2).
Given $\hat{p}$ as in (3.2), an initial $(t,x)\in(\mathbb{T},\mathbb{S}_{t})$, and $\vec{\alpha}\in\vec{\cal A}$, player induces a distribution $\mathbb{P}^{t,x,\vec{\alpha}}:=\mathbb{P}^{\hat{p};t,x,\vec{\alpha}}$ for the canonical process as usual; for all $t\leq s$ and $(\tilde{x},y)\in(\mathbb{S}_{s},\mathbb{S}_{s+1})$, initial condition is $\mathbb{P}^{t,x,\vec{\alpha}}(X_{t}=x)=1$ and
$\mathbb{P}^{t,x,\vec{\alpha}}(X_{s+1}=y|X_{s}=\tilde{x})=\hat{p}(s,\tilde{x},\vec{\alpha}(s,\tilde{x});y).$ (3.3)
@@
(i) We assume that players model the others potential controls depending on their own control. However, one might model that this depends on the path of states of the players, or any other observables are legitimate as long as the cost (3.6) is well-defined.
@@
(ii) For competing players, a sophisticated player might have an estimate on how their actions could be exploitable in order to deceive an opponent, deviating their $\hat{\Gamma}$ from their actual planning. To not only compete with but also cooperate with other players, they may need to generate reliable estimates of the actions of others. In the two-player game discussed in Section 3.1, because the costs to the players depend on each other’s states, omitting this aspect from the player model won’t accurately capture the observed dynamics.
An important notion to introduce is the value of a player. As future observations are ranked by some associated rewards, value function captures a qualitative information about the future rewards for a given strategy. Now, we introduce transition costs and state values: [^3] We use cost and value interchangeably. In the case of scalar objectives as in this work, distinction is more pronounced. However, for multi-objective frameworks, there is typically no binary choice, but rather a continuum of choices.
$\hat{F}:\hat{\Omega}\times\mathbb{T}\times\Omega\times\vec{\mathbb{A}}\to\mathbb{R},\qquad\hat{\phi}:\hat{\Omega}\times\mathbb{T}\times\Omega\to\mathbb{R},\ \$(3.5)
[…]
Now, given $(\hat{T},\hat{p},\hat{\Gamma},\hat{F},\hat{\phi})\in\mathscr{M}_{T}\times\mathscr{M}_{p}\times\mathscr{M}_{\Gamma}\times\mathscr{M}_{F}\times\mathscr{M}_{\phi}$, the value of player becomes
$J(t,x;\alpha)=\int_{\vec{\cal A}}J(t,x;\vec{\alpha})\hat{\Gamma}_{t}(\alpha;d\vec{\alpha}),\ \ \text{where denoting}\ \ \mathbb{E}^{t,x,\vec{\alpha}}=\mathbb{E}^{\mathbb{P}^{t,x,\vec{\alpha}}},$ (3.6)
$J(t,x;\vec{\alpha})=\mathbb{E}^{t,x,\vec{\alpha}}\Big[\hat{\phi}(t+\hat{T},X_{t+\hat{T}})+\sum_{s=t}^{t+\hat{T}-1}\hat{F}(s,X_{s},\vec{\alpha}(s,X_{s}))\Big],\quad$
@@
which is a random variable on $\hat{\Omega}$. Set $\mathscr{M}_{J}^{i}$ as the space of all such functions $(\hat{\Omega}\times\mathbb{T}\times\Omega\times{\cal A}^{i}\to\mathbb{R})$. We point out that requiring a random variable instead of a distribution allows us to refer to samples.
@@
We remark again that a general abstract setting might be a simplification, and there might be many layers of various estimations, such as objects and their relations, before a player actually constructs its value estimate. Indeed, it may be the case for every other estimate too. Transition probabilities might be estimated from simple frequency analysis, or could be modeled by large attention architectures. In fact, a truly complex player, such as a human, would not only adapt existing estimates and behaviors, but also develop new types of estimates and behaviors over time. Just as a child starts with limited capacities and gradually acquires a rich repertoire, such dynamic structural growth might indeed be essential for modeling higher-order intelligence. While this remains far beyond current work, it represents a compelling direction for future research.
Let us recall the game of chess, which serves as an excellent example to keep in mind throughout this work. In chess, $\hat{p}$ yields deterministic transitions. However, a player does not know what actions the opponent will take within $\{t,\dots,t+\hat{T}\}$, and beyond that, it is unclear what the transition costs of actions or the value of being in a particular state at $t+\hat{T}$ might be. These are all crucial components for a player to learn. Notably, the heuristic values assigned to pieces are designed to guide players in learning $\hat{F}$ and $\hat{\phi}$. While simplistic, these heuristics serve as an initial guide. Moreover, as we have mentioned, knowledge about the opponent can alter the values of strategies, which is captured in (3.6) abstractly. Let us also emphasize that the player’s horizon may depend significantly on the current state. Towards the endgame, for instance, a well-trained chess player might be able to estimate many steps ahead, whereas this ability may be considerably more limited during the middle stages of the game.
@@
And then, to incorporate the potential difficulty and uncertainty in identifying the optimal control, we introduce the next policy estimation;
$\hat{\pi}:\hat{\Omega}\times\mathbb{T}\times\Omega\to{\cal A}^{i}\qquad\text{where},$ (3.7)
$\hat{\pi}(\hat{\omega},t,x)\in{\cal A}^{t,x;i},\quad\forall(\hat{\omega},t,x)\in\hat{\Omega}\times\mathbb{T}\times\mathbb{S}_{t}$
Here, at $(\hat{\omega},t,x)$, $\hat{\pi}$ approximates the potential optimal controls for $J(\hat{\omega},t,x,\cdot)$, which will be dictated by the equilibrium condition below. We remark that $\hat{\Omega}$ might have a component for both value and policy, which we don’t keep track explicitly. For example, we might have a collection of events where the value is determined, whereas $\hat{\pi}$ is still a random variable. Set $\mathscr{M}_{\pi}$ as the set of functions as in (3.7).
@@
Now, even when optimal control can be solved exactly, uncertainty over the value will naturally induce a probability distribution over controls. First, introduce the space of behavior $\mathscr{M}_{\Upsilon}$ as functions of the form
$\hat{\Upsilon}:\hat{\Omega}\times\mathbb{T}\times\Omega\to\mathbb{A}^{i}$ (3.8)
Then, we introduce a behavior in a straightforward manner from the current estimates as follow;
${\Upsilon}:\Pi_{\varphi}\mathscr{M}_{\varphi}\to\mathscr{M}_{\Upsilon}\quad\text{where, }$ (3.9)
${\Upsilon}(\hat{\pi})=((\hat{\omega},t,x)\mapsto\hat{\pi}(\hat{\omega},t,x)(t,x))\quad$
@@
Here, behavior is choosing the immediate action assigned by the policy, and reflects the randomness induced by both the value and policy. This behavior is typical for the learning or playing phase and is, in essence, similar to Thompson sampling adapted to our framework. During a competition phase, one might choose a different $\Upsilon$} as deterministic, selecting the action corresponding to the mode of them. In particular situations, such as performing surgery, it is not only wrong but also unethical to forgo the most likely action and instead select one at random. Moreover, a player may adjust its behavior to occasionally select unlikely actions, exploring states that are disadvantageous or even entirely unseen when facing a weak opponent.
@@
It is important to motivate the role of randomness in value, which then induces a probability distribution over actions by (3.9). Recall that in sufficiently complex settings, such as chess, values are inherently unknown and must be learned through significant effort. That is, the randomness of the value models what is unknown to the player. One key role of randomness in value is to allow players to explore systematically. If the player is not satisfied with the current value estimates, it is natural to shift the estimates for unexplored controls, or their outcomes, toward higher values, in anticipation that they may achieve better results than current estimates. This approach naturally leads the player to search for controls yielding more satisfactory outcomes. We will demonstrate a toy version for the two-player game in Section 3.1. This approach aligns with the common intuition that a better understanding of values should lead to less uncertain strategies.
Now that we have introduced the spaces of estimations and behavior, let us turn to players as in Definition 1. Observations may come from real-world experience, or, in the mean-field regime, players can generate observations by assuming that every other player is identical. For multiple players, we define the observations as mappings of the form
${\cal O}^{i}:\Omega^{u}\times\vec{\mathscr{A}}\times\mathbb{N}\to\mathscr{E}^{i}$ (3.10)
satisfying consistency as in Definition 2. We then set $\vec{\cal O}=({\cal O}^{1},{\cal O}^{2},\dots)$.
@@
Next, we formally acknowledge the existence of learning algorithms. We say a collection of functions $\mathfrak{L}_{\varphi}^{i}$ for $\varphi\in\{T,p,\Gamma,F,\phi,\pi\}$ is the learning algorithm of player $i$. Recall that $\mathscr{M}_{T},\mathscr{M}_{p},\mathscr{M}_{\Gamma},\mathscr{M}_{F},\mathscr{M}_{\phi},\mathscr{M}_{\pi}$ respectively denote the spaces of estimations as in (3.1), (3.2), (3.4), (3.5), and (3.7) respectively. Then, learning algorithms are in general of the form
@@
$\mathfrak{L}_{\varphi}^{i}:\mathscr{E}^{i}\times\Pi_{\tilde{\varphi}}\mathscr{M}_{\tilde{\varphi}}\times\mathscr{M}_{\Upsilon}\to\mathscr{M}_{\varphi},\quad\forall\varphi\in\{T,p,\Gamma,F,\phi,\pi\}.$ (3.11)
Let us remark that, although it was not necessary, we introduced $\Upsilon$} explicitly given the other estimations. In its given form, it is a value-based approach in reinforcement learning (see Section 5). However, learning algorithms in (LABEL:eqn:learningalgorithms) are crucial, and we do not attempt to simplify them. For example $p$ and $\Gamma$ might be defined directly from observations by keeping frequency statistics. Although this allows them to be more trackable, they are limited to simple settings. Our motivation in this work is to emphasize their inherent complexity instead. They must be subject to evaluations of their respective objectives, such as in Definition 5.
Lastly, to introduce the estimates and the planned behavior of players, we denote the priors as
$^{0}\vec{I}\in\vec{\mathscr{A}},\quad\text{and}\quad{}^{0}\vec{\mathfrak{L}}_{\varphi}\in\vec{\mathscr{M}}_{\varphi},\quad\forall\varphi\in\{T,p,\Gamma,F,\phi,\pi\}.$
This sets $^{0}\vec{\Upsilon}$ as in (3.9). Then,
${}^{n}{\Upsilon}^{i}$ $={\Upsilon}({}^{n}\mathfrak{L}_{\varphi}^{i})\in\mathscr{M}_{\Upsilon},\quad\forall i\in\mathbb{N}_{0}$ (3.12)
@@
${}^{n}\mathfrak{L}_{\varphi}^{i}$ $=\mathfrak{L}_{\varphi}({}^{n}{\cal O}^{i},\ {}^{n-1}\mathfrak{L}_{\varphi}^{i},\ {}^{n-1}{\Upsilon}^{i})\in\mathscr{M}_{\varphi},\quad\forall\varphi\in\{T,p,\Gamma,F,\phi,\pi\},i\in\mathbb{N}_{0}$
${}^{n}\vec{\cal O}$ $={\cal O}(\omega^{u},{}^{n-1}I,n)\in\vec{\mathscr{E}},$
${}^{n}\vec{I}$ $=({}^{0}\vec{I},{}^{1}\vec{\Upsilon}(\hat{\omega},t,x),\dots,{}^{n}\vec{\Upsilon}(\hat{\omega},t,x))\in\vec{\mathscr{A}}.$
Also, $^{n}(\mathfrak{L}_{T},\mathfrak{L}_{p},\mathfrak{L}_{\Gamma},\mathfrak{L}_{F},\mathfrak{L}_{\phi})^{i}$ defines $^{n}J^{i}(t,x;\alpha)$ as in (3.6), and we may use the notation $^{n}\hat{\varphi}^{i}:={}^{n}\mathfrak{L}_{\varphi}^{i}$ if convenient.
@@
We say that players $(\vec{\cal O},\vec{\mathfrak{L}}_{\varphi},\vec{\Upsilon})$ admit $^{*}\vec{\Upsilon}\in\vec{\mathscr{M}}_{\Upsilon}$ as an $(\varepsilon,r,\delta)$-uncertain equilibrium under metrics $d^{i}$ on $\mathscr{M}_{\Upsilon}^{i}$, if for any prior $^{0}\vec{I}\in\vec{\mathscr{A}}$,